# Cosmic-Ray Feedback from Supernovae in a Stratified Interstellar Medium

### v1

@@ 3.3 Saturated State
Second, we can examine the cosmic-ray transport in the midplane compared to the transport in the extraplanar gas. In the top row of Figure 14, we show the cosmic-ray pressure-gas number density phase space. The histograms are averaged over the steady state time frame of $t\in\left[300,400\right]\,\mathrm{Myr}$. These figures show a clear differentiation in cosmic-ray transport between the dense gas and diffuse gas. In the dense gas, cosmic rays rapidly diffuse, ending up almost uniformly distributed across orders of magnitude in gas density. Once the cosmic rays escape the dense gas, they enter the extraplanar gas which has a lower density, and high speeds, allowing advective and streaming transport to dominate. In both our simulations, the streaming transport is dominant. This streaming transport is apparent in the histograms, which do not follow the $\gamma_{c}=4/3$ polytropic law $(P_{c}\propto n^{\gamma_{c}})$ for cosmic rays advected with the thermal gas. Instead, it follows a $\gamma_{c}=2/3$ index, which is expected for a steady state system with streaming transport (Wiener et al., 2019). This result is unexpected given advective and diffusive transport are also included in the simulation. Likely, since the magnetic field is predominantly vertical and aligned with the cosmic-ray pressure gradient, the streaming becomes dominant. In the CRInj simulation, there is also a spike just below $\log_{10}(n/(1\mathrm{cm}^{-3}))=0$. This spike comes from averaging over all the cosmic-ray injections which occur during the steady state time frame. Those injections are often in gas which is heated (moved to lower densities) by thermal feedback, and where cosmic rays propagate along the magnetic field.
@@
Figure 14 also shows the gas pressure-gas number density phase space (middle row) and the gas temperature-gas number density phase space (bottom row). The histograms are averaged over the steady state time period. For the most part, the profiles in both simulations follow the thermal equilibrium curve of $\Gamma=n\Lambda$ shown in Figure 2 and plotted as a dashed green line in middle row plots. However, the CRInj simulation’s histogram departs from this curve at low gas densities, and this low density gas is also at a higher temperature in the temperature-density phase space. This heating comes from the streaming transport of cosmic rays, which takes over in the diffuse gas. At the lower density, the net cooling time increases and is unable to offset cosmic-ray heating, leading to increased temperature. While we have not done any simulations in which streaming is present but cosmic-ray heating is removed from the energy equation, the relative paucity of low density, high temperature in the NoCRInj model demonstrates that this is an important effect.
@@
After this examination of the steady state, a clear picture of the interplay between magnetic field orientation and cosmic-ray feedback emerges. The shift to a predominantly vertical magnetic field outside the midplane opens the door for more effective cosmic-ray feedback. While injected cosmic rays diffuse in the midplane, they stream and provide a large pressure gradient outside the midplane, producing a faster wind compared to the simulation without cosmic-ray injections. The streaming of cosmic rays also causes them to heat the thermal gas, producing hotter gas.
@@ 3.4 Observational Comparison
Figure 15: Average vertical profiles from the CRInj simulation, fitted with two independent scale heights. The red shaded regions show the standard deviation of the vertical profiles during the saturated time frame, and the solid red lines show the median profile. The dashed green lines show the initial hydrostatic equilibrium, and the dashed black lines show the best fit to each profile. The profiles are produced by first removing all cold gas with $T<500\,\mathrm{K}$
Figure 16: Ratio of gas scale heights between the CRInj simulation and the NoCRInj simulation. Black lines show the ratio of gas density scale heights, and red lines show the ratio of gas pressure scale heights. Solid lines show the outer scale height (for the lower amplitude, larger scale height component of the fits), and dashed lines show the inner scale height (for the higher amplitude, smaller scale height component of the fits).
@@
Figure 15 shows the average vertical profiles and our median fit to those profiles during the saturated, steady state. The averaged data are plotted within a shaded region, which shows the total variation over the saturated time frame. We removed the cold and thermally unstable gas disk in the midplane, which is not well-resolved, by using only gas cells with a temperature $T>3000\,\mathrm{K}$ for the $xy$-plane averaging. We use a linear combination of the base hydrostatic equilibrium profile (see Section 2.1):
@@
At each $1\,\mathrm{Myr}$ time dump, this gives us a pair of scale heights $H_{\mathrm{inner}}$ $H_{\mathrm{outer}}$ and amplitudes $A_{\mathrm{inner}}$ $A_{\mathrm{outer}}$. The inner amplitude is always larger than the outer amplitude, and the inner scale height is always smaller than the outer scale height. We chose the above profile because of our gravitational acceleration profile. The fitted profiles are pictured for gas density and gas pressure in Figure 15, with a central peak near the midplane and large wings oustide the midplane.
@@
While the fitted scale heights we can be related to an exponential scale height for $z\geq H_{*}$, we caution readers against a direct comparison with observed scale heights in the Milky Way. This caution is only because of the less realistic gravitational acceleration profile we use (see Section 2.1). Instead, we encourage a focus on the structure which our simulations produce: a thinner, dense disk and a thicker disk inflated by non-thermal pressures. A two scale height model matches with the modern understanding of the structure of the Milky Way’s thermal gas (McClure-Griffiths et al., 2023; Rybarczyk et al., 2024).
@@
For the non-thermal pressures, there is a different story. The cosmic-ray pressure profile in the third row of Figure 15 looks significantly different and cannot be fit by the hydrostatic equilibrium profile in Equation 22. The magnetic energy density profile at the bottom still has an inner profile and an outer profile, each following Equation 22. However, the outer profile only takes over at $z\approx\pm 1.5\,\mathrm{kpc}$ instead of at $z\approx\pm 0.4\,\mathrm{kpc}$ like the gas density and pressure. Figure 15 also illustrates that the gas density, magnetic field, and cosmic rays do not trace each other well.
@@
In Figure 16, we show the time evolution of scale heights of gas density and magnetic energy density over the saturated state for both simulations. We plot these quantities because they are measurable (e.g. gas density via HI emission and absorption, and magnetic energy density through synchrotron radiation). We only plot the inner profile for the magnetic energy density because the values we measure for the outer scale height $(\gtrsim 2\,\mathrm{kpc})$ are untrustworthy. Those large values for outer scale height of the magnetic energy density are close to the simulation’s box size and the region is very close to the vertical outflow boundaries. The only point to be taken away from the outer scale height of the magnetic energy is that it flattens out (see bottom plot of Figure 15) due to the increase in vertical magnetic field strength at large $z$ (see Figure 13).
In the top plot of Figure 16, we see that the scale heights are relatively steady during the saturated state. The inner and outer density scale heights differ by a large factor (×5)($\times$ 5)( × 5 ), and the magnetic energy density’s inner scale height falls in between the density scale height values. While the inner density scale height and inner magnetic energy density scale height change little when comparing the two simulations, the outer density scale height increases by over $100\,\mathrm{pc}$. This inflation of the outer disk by cosmic-ray injections is more quantitatively shown in the bottom plot of Figure 16, where we take the ratio of the scale heights from each simulation. The inner gas density and magnetic energy density scale height ratio stays near $1$ for most of the simulation, showing the cosmic-ray injections have a minimal impact inside the midplane region $|z|\leq 400\,\mathrm{pc}$. However, the cosmic-ray injections inflate the outer gas density scale height by between $25\%$ and $50\%$.
@@
There is another observable quantity to consider: $\gamma$-ray luminosity. This luminosity is proportional to the cosmic-ray energy density, providing a useful tool for probing the cosmic-ray content of galaxies other than the Milky Way. Cosmic rays can produce pions when they collide with non-relativistic nuclei. Some of those pions, about a third, emit $\gamma$-rays as they decay. In Figure 17, we show the evolution of the total energy lost in all hadronic interactions for our simulations. We normalize this by the cosmic-ray energy injection rate in the CRInj simulation, following the definition of the $F_{\mathrm{cal}}=L_{\pi}/\dot{E}_{\mathrm{cr}}$ parameter in Lacki et al. (2011). We calculate the hadronic loss rate $L_{\pi}$ using the formula for the hadronic loss rate per volume $\Gamma_{h}$ in Equation 12 of Guo & Oh (2008):
@@
In the bottom plot of Figure 17, we show $F_{\mathrm{cal}}$ for both simulations (note that this is not the best measure for the NoCRInj simulation, which should really have an undefined $F_{\mathrm{cal}}$). Since most of the gas density is in the midplane and $L_{\mathrm{\pi}}$ is proportional to gas density (see Equation 23), most of the hadronic losses occur in the midplane. Initially, the losses are large because of the initial hydrostatic equilibrium we forced on the system (see Section 2.1). After $t\approx 250\,\mathrm{Myr}$, the hadronic losses in the NoCRInj simulation vanish much of the initial cosmic-ray background has escaped out of the top and bottom of the simulation. The emission outside the midplane in the CRInj simulation also vanishes. However, in the saturated, steady state we examined in Section 3.3, the only hadronic losses level off in the midplane of the CRInj simulation.
@@
To get to a physically realistic and comparable $F_{\mathrm{cal}}$ parameter, we need to get rid of the initial background cosmic-ray pressure the simulation started with. For an approximation of this, we subtract the NoCRInj hadronic loss rate off of the CRInj hadronic loss rate. This net calorimetric fraction is plotted in the top of Figure 17 for the midplane, outside the midplane, and their combination. Initially, as cosmic-ray injections are dumped into the simulation volume, the $F_{\mathrm{cal}}$ increases at a constant rate. Then, at $100\,\mathrm{Myr}$, when the Parker instability begins producing a strong vertical magnetic field, (see Figure 7), the net $F_{\mathrm{cal}}$ decreases nearly as rapidly as it rose. With the vertical magnetic field, the cosmic-ray injections are able to escape the simulation faster, producing less hadronic interactions.
@@
As the simulation saturates, we see a steady growth of the net $F_{\mathrm{cal}}$ parameter. However, its growth is at a lower rate than early in the simulation, despite our cosmic-ray injection rate being constant (the injections are stochastic in time, but the average rate is constant - see Section 2.3). Since there is still some vertical magnetic fields at these times, there is more escape than at the beginning of the simulation. But that escape is still not enough to agree with observations from Lacki et al. (2011). The only time our calorimetric fraction gets close to the observational values (basically, under $0.5$) is at the saturation of the Parker instability, near $t\sim 200\,\mathrm{Myr}$.
@@
Figure 17: Bottom: Fraction of cosmic-ray energy injected which would be lost to hadronic interactions, following the definition of $F_{\mathrm{cal}}$ in Lacki et al. (2011). We plot the results for both simulations, with solid lines showing the CRInj simulation and dashed lines showing the NoCRInj simulations. We separate the emission from the midplane $|z|\leq H_{*}$, shown in purple, from the emission outside of the midplane, shown in green. Top: We calculate the Net $F_{\mathrm{cal}}$ by subtracting the NoCRInj simulation’s $F_{\mathrm{cal}}$ from the CRInj simulation’s $F_{\mathrm{cal}}$. This subtraction leaves only the $F_{\mathrm{cal}}$ associated with our cosmic-ray injections. For comparison, the blue shaded region shows observational estimates of $F_{\mathrm{cal}}$ in starbursting galaxies NGC 253 and M82 (Lacki et al., 2011).
@@ 4 Discussion
Returning to the original question, we find including cosmic-ray injections did make a difference. Changing the stellar feedback by putting $10\%$ of supernova energy into cosmic rays led to a stronger outflow. It also produced a steady outflow, reaching from just $500\,\mathrm{pc}$ above and below the midplane out to $2.4\,\mathrm{kpc}$. This steady outflow was driven by a cosmic-ray pressure gradient, which was fueled by injected cosmic rays streaming along vertical magnetic field lines.
@@
However, there were also similarities between the two simulations. The cosmic-ray injections had a minimal impact on the vertical density structure (see Figure 8). The vertical density profiles in the steady state were similar, with the only difference being an inflation of the thick disk scale height by a factor of $\gtrsim 25\%$ (see Figure 16).
@@
Our results also suggest some solutions to other questions and problems. The CRInj simulation illustrated that the orientation of the magnetic field can have a significant impact on cosmic-ray transport. The change in transport has implications for the cosmic-ray calorimeter problem and the vertical stability of the ISM. Before we address these in Sections 4.1 & 4.2, note that these implications are not isolated to the Parker instability. The change in cosmic-ray transport we see is due to the vertical magnetic field outside of the midplane. This vertical orientation could result from the Parker instability, as it did in our simulations, or from other mechanisms.
4.1 Cosmic-Ray Calorimeter Problem

### v2

@@ 3.3 Saturated State
Second, we can examine the cosmic-ray transport in the midplane compared to the transport in the extraplanar gas. In the top row of Figure 15, we show the cosmic-ray pressure-gas number density phase space. The histograms are averaged over the steady state time frame of $t\in\left[260,400\right]\,\mathrm{Myr}$. These figures show a clear differentiation in cosmic-ray transport between the dense gas and diffuse gas. In the dense gas, cosmic rays rapidly diffuse, ending up almost uniformly distributed across orders of magnitude in gas density. Once the cosmic rays escape the dense gas, they enter the extraplanar gas which has a lower density, and high speeds, allowing advective and streaming transport to dominate. This transition in transport matches what has been seen in previous tallbox simulations with post-processing of cosmic-ray transport, highlighting a transition to advective transport (Armillotta et al., 2021, 2024).
In both our simulations, the streaming transport is dominant over advection. This streaming transport is apparent in the histograms, which do not follow the $\gamma_{c}=4/3$ polytropic law $(P_{c}\propto n^{\gamma_{c}})$ for cosmic rays advected with the thermal gas. Instead, it follows a $\gamma_{c}=2/3$ index, which is expected for a steady state system with streaming transport (Wiener et al., 2019). This result is unexpected given our system is not in a perfect steady state and other transport mechanisms (advection and diffusion) are also included in the simulation. Likely, since the magnetic field is predominantly vertical and aligned with the cosmic-ray pressure gradient, the streaming becomes dominant. Additionally, the simulations have a slightly stronger magnetic field than other works, and the turbulence is weak. These changes amplify streaming as a transport mechanism over advection.
@@
Figure 15 also shows the gas pressure-gas number density phase space (middle row) and the gas temperature-gas number density phase space (bottom row). The histograms are averaged over the steady state time period. For the most part, the profiles in both simulations follow the thermal equilibrium curve of $\Gamma=n\Lambda$ shown in Figure 2 and plotted as a dashed green line in middle row plots. However, the CRInj simulation’s histogram departs from this curve at low gas densities, and this low density gas is also at a higher temperature in the temperature-density phase space. This heating comes from the streaming transport of cosmic rays, which takes over in the diffuse gas. At the lower density, the net cooling time increases and is unable to offset cosmic-ray heating, leading to increased temperature. While we have not done any simulations in which streaming is present but cosmic-ray heating is removed from the energy equation, the relative paucity of low density, high temperature in the CRBkg model demonstrates that this is an important effect.
@@
After this examination of the steady state, a clear picture of the interplay between magnetic field orientation and cosmic-ray feedback emerges. The shift to a predominantly vertical magnetic field outside the midplane opens the door for a cosmic-ray driven outflow. While injected cosmic rays diffuse in the midplane, they stream and provide a large pressure gradient outside the midplane, producing a faster wind compared to the simulation without cosmic-ray injections. The streaming of cosmic rays also causes them to heat the thermal gas, producing a hotter outflow.
@@ 3.4 Observational Comparison
Figure 16: Average vertical profiles from the CRInj simulation, fitted with two independent scale heights. The red shaded regions show the standard deviation of the vertical profiles during the saturated time frame, and the solid red lines show the median profile. The dashed green lines show the initial hydrostatic equilibrium, and the dashed black lines show the best fit to each profile. The profiles are produced by first removing all cold and unstable gas with $T<3000\,\mathrm{K}$
Figure 17: Gas scale heights in CRInj simulation and CRBkg simulation. Black lines show the inner warm gas density scale heights, and red lines show the outer warm gas density scale heights. The green lines show the scale height for the magnetic energy density, and blue lines show the scale height of the thin, cold disk. Solid lines show CRInj simulation, and dashed lines show the CRBkg simulation.
@@
Figure 16 shows the average vertical profiles and our median fit to those profiles during the saturated, steady state. The averaged data are plotted within a shaded region, which shows the total variation over the saturated time frame. We removed the cold and thermally unstable gas disk in the midplane to isolate the impact on warm and hot gas, by using only gas cells with a temperature $T>4000\,\mathrm{K}$ for the $xy$-plane averaging. We use a linear combination of the base hydrostatic equilibrium profile (see Section 2.1):
@@
At each $1\,\mathrm{Myr}$ time dump, this gives us a pair of scale heights $H_{\mathrm{inner}}$ $H_{\mathrm{outer}}$ and amplitudes $A_{\mathrm{inner}}$ $A_{\mathrm{outer}}$. We chose the above profile because it is the equilibrium profile under the influence of our chosen gravitational acceleration (Equation 12). The fitted profiles are pictured for gas density and gas pressure in Figure 16, with a central peak near the midplane and large wings outside the midplane.
@@
While the fitted scale heights can be related to an exponential scale height for $z\geq H_{*}$, we do not expect them to always match observed scale heights in the Milky Way. This expectation is because the gravitational acceleration profile we use is a simplification of the true acceleration profile. Instead, we encourage a focus on the structure which our simulations produce: a thinner, dense disk and a thicker disk inflated by non-thermal pressures. A two scale height model matches with the modern understanding of the structure of the Milky Way’s thermal gas (McClure-Griffiths et al., 2023; Rybarczyk et al., 2024).
@@
For the non-thermal pressure, there is a different story. The cosmic-ray pressure profile in the third row of Figure 16 looks significantly different and cannot be fit by the hydrostatic equilibrium profile in Equation 22. The magnetic energy density profile in the fourth plot still has an inner profile and an outer profile, each following Equation 22. However, the outer profile only takes over at $z\approx\pm 1.5\,\mathrm{kpc}$ instead of at $z\approx\pm 0.4\,\mathrm{kpc}$ like the gas density and pressure. A similar structure also appears for the turbulent pressure in the bottom plot. Figure 16 also illustrates that the gas density, magnetic field, and cosmic rays do not trace each other well.
@@
In Figure 17, we show the time evolution of scale heights of gas density and magnetic energy density over the saturated state for both simulations. We also include the scale height for the cold gas, which we find by fitting a single component with the form of Equation 22. We plot these quantities because they are measurable (e.g. gas density via HI emission and absorption, and magnetic energy density through synchrotron radiation). We only plot the inner profile for the magnetic energy density because the values we measure for the outer scale height $(\gtrsim 2\,\mathrm{kpc})$ are untrustworthy. The outer scale height of the magnetic energy density we find is close to the simulation’s vertical box size $(4.8\mathrm{kpc})$, and the outer scale height only takes over above $1.5\mathrm{kpc}$ (see Figure 16). The only point to be taken away from the outer scale height of the magnetic energy is that it flattens out due predominantly vertical magnetic field at large $z$ (see Figure 14).
In Figure 17, it is apparent that the inner scale height of the warm gas, the inner scale height of the magnetic energy density, and the scale height of cold gas, are relatively constant during the steady state time frame $(t>260\,\mathrm{Myr})$. The outer scale height of the warm gas varies significantly in both simulations. The inner and outer warm gas scale heights differ by a large factor (×5)($\times$ 5)( × 5 ), and the magnetic energy density’s inner scale height falls in between the two warm gas components. On average, the CRInj has a slightly larger outer warm gas scale height component and larger magnetic scale height. But the simulations are in overall agreement with each other, and both illustrate the need for two components to fit the warm gas’ structure.
While the cold gas component initially has a scale height of $\lesssim 50\,\mathrm{pc}$, it eventually decays to a value equal to the resolution of our simulation $\sim 5\,\mathrm{pc}$. Without a more complete feedback prescription (i.e. increase the supernovae rate as more cold gas forms), there is nothing stopping the cold gas from falling to $z=0$ (see Section 4.3) .
@@
Another observable quantity to consider is $\gamma$-ray luminosity due to hadronic collisions of cosmic rays with thermal particles. These collisions produce a cascade of reactions which includes the emission of $\gamma$-rays. This luminosity is proportional to the cosmic-ray energy density, providing a useful tool for probing the cosmic-ray content of galaxies other than the Milky Way (Guo & Oh, 2008). In Figure 18, we show the evolution of the total energy lost in all hadronic interactions for our simulations, not just the energy emitted in $\gamma$-rays. We normalize this by the cosmic-ray energy injection rate in the CRInj simulation, following the definition of the $F_{\mathrm{cal}}=L_{\pi}/\dot{E}_{\mathrm{cr}}$ parameter in Lacki et al. (2011). We calculate the hadronic loss rate $L_{\pi}$ using the formula for the hadronic loss rate per volume $\Gamma_{h}$ in Equation 12 of Guo & Oh (2008):
@@
In the bottom plot of Figure 18, we show $F_{\mathrm{cal}}$ for both simulations. The value can be larger than one in our simulations because we do not actually include the loss rate in our simulations. Additionally, these values are not the true measure in the case of the CRBkg simulation which actually has $\dot{E}_{\mathrm{cr,inj}}=0$.
@@
Since most of the gas density is in the midplane and $L_{\mathrm{\pi}}$ is proportional to gas density (see Equation 23), most of the hadronic losses occur in the midplane. Initially, the losses are large because of the initial hydrostatic equilibrium we forced on the system (see Section 2.1). After $t\approx 250\,\mathrm{Myr}$, the hadronic losses in the CRBkg simulation vanish because much of the initial cosmic-ray background has escaped out of the top and bottom of the simulation. The emission outside the midplane in the CRInj simulation also vanishes. However, in the saturated, steady state we examined in Section 3.3, the hadronic losses level off to a steady value in the midplane of the CRInj simulation.
@@
To get to a physically realistic and comparable $F_{\mathrm{cal}}$ parameter, we need to get rid of the initial background cosmic-ray pressure. For an approximation of this, we subtract the CRBkg hadronic loss rate off of the CRInj hadronic loss rate. This net calorimetric fraction is plotted in the top of Figure 18 for the midplane, outside the midplane, and their combination. Initially, as cosmic-ray injections are dumped into the simulation volume, the net $F_{\mathrm{cal}}$ increases at a constant rate. Then, at $100\,\mathrm{Myr}$, when the Parker instability begins producing a strong vertical magnetic field, (see Figures 6 and 8), the net $F_{\mathrm{cal}}$ decreases nearly as rapidly as it rose. With the vertical magnetic field, the cosmic-ray injections are able to escape the simulation faster, producing less hadronic interactions.
@@
As the simulation saturates, we see a steady growth of the net $F_{\mathrm{cal}}$ parameter. However, its growth is at a lower rate than early in the simulation, despite our cosmic-ray injection rate being constant (the injections are stochastic in time, but the average rate is constant - see Section 2.3). Since there is still some vertical magnetic fields at these times, there is more escape than at the beginning of the simulation. But that escape is still not enough to agree with observations from Lacki et al. (2011). The only time our calorimetric fraction gets to the observational values is at the nonlinear saturation of the Parker instability, near $t\sim 180\,\mathrm{Myr}$ when the vertical magnetic field strength is largest (see Figure 3).
@@
Figure 18: Bottom: Fraction of cosmic-ray energy injected which would be lost to hadronic interactions, following the definition of $F_{\mathrm{cal}}$ in Lacki et al. (2011). We plot the results for both simulations, with solid lines showing the CRInj simulation and dashed lines showing the CRBkg simulations. We separate the emission from the midplane $|z|\leq H_{*}$, shown in purple, from the emission outside of the midplane, shown in green. Top: We calculate the Net $F_{\mathrm{cal}}$ by subtracting the CRBkg simulation’s $F_{\mathrm{cal}}$ from the CRInj simulation’s $F_{\mathrm{cal}}$. This subtraction leaves only the $F_{\mathrm{cal}}$ associated with our cosmic-ray injections. For comparison, the blue shaded region shows observational estimates of $F_{\mathrm{cal}}$ in starbursting galaxies NGC 253 and M82 (Lacki et al., 2011).
Figure 19: Top plot: Net $F_{\mathrm{cal}}$ as calculated in Figure 18, in both our low resolution and high resolution simulations. Bottom plot: Slices of gas density and magnetic fields in the low resolution and high resolution simulations at $t=100\mathrm{Myr}$ and $t=180\mathrm{Myr}$. The vertical magnetic field facilitates the escape of cosmic rays from the midplane, which decreases the $\gamma$-ray luminosity and the net $F_{\mathrm{cal}}$.
This minimum point in evolution of $F_{\mathrm{cal}}$ also depends on resolution. In Figure 19, we show the evolution of the net calorimetric fraction over time for a $10\,\mathrm{pc}$ resolution simulation and for our fiducial $5\,\mathrm{pc}$ resolution CRInj simulation. Both simulations follow a similar evolution over time, although the higher resolution simulation shows a steeper decrease in net $F_{\mathrm{cal}}$ during the saturation of the Parker instability.
The bottom part of Figure 19 shows how $F_{\mathrm{call}}$ decreases: the magnetic field becomes vertical between $t=100\,\mathrm{Myr}$ and $t=180\,\mathrm{Myr}$. These field lines allow cosmic rays to escape the simulation, decreasing the $\gamma$-ray luminosity. After the pictured slices, the vertical field lines begin to numerically reconnect. This numerical reconnection is not physical, but it does disconnect the cosmic rays in the midplane from the outflow boundary conditions in the $\hat{z}$ direction. This disconnect leads to an overestimation of $\gamma$-ray luminosity.
@@ 4 Discussion
Returning to the original goal of understanding how cosmic ray injections impact stellar feedback, we have seen their primary effect is in accelerating gas at $|z|\gtrsim 0.5\,\mathrm{kpc}$, building the base behind a galactic wind. This steady outflow lasted for $\gtrsim 100\,\mathrm{Myr}$ and was primarily driven by the cosmic-ray pressure gradient. That gradient was sourced by cosmic rays streaming along vertical magnetic field lines and it was amplified when cosmic ray injections were included. This result is similar to what has been seen in galaxy scale simulations like Farber et al. (2018) and Thomas et al. (2024), where cosmic rays drive a low density outflow and dense gas is left behind. In fact, our result could form the base of these large scale outflows, or source an outflow along an open magnetic flux tube (Breitschwerdt et al., 1991; Shimoda & Inutsuka, 2022).
@@
However, there were also similarities between the two simulations. The cosmic-ray injections had a minimal impact on the vertical density structure (see Figure 9). The vertical density profiles of each simulation in the steady state were similar, with the only difference being in the evolution of the scale height of warm gas outside the midplane.
Additionally, the Parker instability developed in both simulations and drove a fast outflow while reorienting the magnetic field. The instability launched a short lived and massive outflow before the simulations settle into a steady, but weaker, outflow. The outflow slowed because the free energy (contained in the gravitational stratification and non-thermal pressures) of the initial conditions was released during the nonlinear saturation of the instability. The end effect of the Parker instability in the simulations was the creation of a vertical magnetic field, which facilitated cosmic ray escape and the formation of cold dense gas.
@@
Our results also suggest some solutions to other questions and problems. The CRInj simulation illustrated that the orientation of the magnetic field has a significant impact on cosmic-ray transport. The change in transport has implications for the cosmic-ray calorimeter problem and the vertical stability of the ISM. Before we address these in Sections 4.1 & 4.2, note that these implications are not isolated to the Parker instability. The change in cosmic-ray transport we see is due to the vertical magnetic field outside of the midplane. This vertical orientation could result from the Parker instability, as it did in our simulations, or from other mechanisms like superbubble chimneys (O’Neill et al., 2024).
4.1 $\gamma$-Ray Bright Galaxy Problem